A firm paid $500,000 for a manufacturing facility. It will produce 45 tons of packing material for each of the first 3 years that will be sold for $4000/ton. For the next 5 years, it will produce 60 tons of material that will be old for $5,000/ ton. If i=15% and the annual operating costs were $100,000, did the firm make a good investment?
Answer #1

annual cash flow for the first 3 years=(45*4000)-100000=80000

annual cash flow for the next 5 years=(60*5000)-100000=200000

NPV of the investment

=-500000+80000/(1+15%)^1+80000/(1+15%)^2+80000/(1+15%)^3+200000/(1+15%)^4+200000/(1+15%)^5+200000/(1+15%)^6+200000/(1+15%)^7+200000/(1+15%)^8

=123477.29
as NPV is positive, so this is good investment

the above is answer..